OpenStudy (anonymous):

Can someone please help me, and show me the steps on how to do this? What is the value of x?

OpenStudy (anonymous):

The sum of angels in a triangle is 180. so \[180 - (50.9 + 33.2) = 180 - 84.1 = 95.9\] The angels in poing G are vertical angles, means that they are even so the other angle in G is also 95.9. The sum of the internal angels in a pentagon (5-sided polygon) is 540. so \[95.9 + 130.4 + 129.6 + 90 = 345.9\] \[x = 540 - 445.9 = 94.1\]

OpenStudy (anonymous):

correction: \[95.9 + 130.4 + 129.6 + 90 = 445.9\]

OpenStudy (anonymous):

Thank You so much I get it now, That's easy! Thanks :)

OpenStudy (anonymous):

No problem.